Q.2 The quickest method to compute the probability that by choosing

by chance 3 pupils, one out of each school, at least one of them wears glasses, is to evaluate the probability that none of them wears glasses. If B is the event that at least one of the 3 pupils wears glasses. 


1
Expert's answer
2021-07-11T17:35:31-0400

We consider three different elementary schools "E, M, S."

The percentage of pupils wearing glasses is 10% in the school "E," 25% in the school "M" and 40% in the school "S."

If "B" is the event that at least one of the 3 pupils wears glasses, then


"P(B)=1-P(\bar{B})"

"P(\bar{B})=\dfrac{100-10}{100}(\dfrac{100-25}{100})(\dfrac{100-40}{100})"

"=\dfrac{81}{200}=0.405"

Then


"P(B)=1-\dfrac{81}{200}=\dfrac{119}{200}=0.595"

The probability that by choosing randomly 3 pupils, one out of each school, at least one of them wears glasses is "\dfrac{119}{200}=0.595."